A $50,000, level-coupon Eurobond has a 6% coupon and matures in ten years. At what price should the bond sell today if the prevailing interest rate is 8% per annum? Round your answer to the nearest dollar.

A) $43,290

B) $29,870

C) $50,000

D) $57,360
Answer #1

Annual coupon=50,000*6%=$3000

Hence price of bond=Annual coupon*Present value of annuity factor(8%,10)+$50,000*Present value of discounting factor(8%,10)

=3000*6.710081399+$50,000*0.463193488

=$43290(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=3000[1-(1.08)^-10]/0.08

=3000*6.710081399

2.Present value of discounting factor=50,000/1.08^10

=50,000*0.463193488